# Pushing the limits of eccentricity in planet-disc interactions

### v1

@@ 2.4 Important diagnostics
[…]
@@
[…]
These quantities can be computed at any instant from the net perturbation maps. Alternatively, pursuing a more elegant approach, one may insert the modal expansion for each perturbed variable, given by equation (3), into equations (13)-(14) and then exploit integral orthogonality conditions to simplify matters (see FR25). Furthermore, we are interested in the orbit-averaged quantities enacted through the operator $\langle\cdot\rangle=(1/T_{\rm p})\oint\cdot dt$, where $T_{\rm p}=2\pi/n$. This averaging simplifies the expression further and we arrive at equations which encapsulate the individual modal contributions to the AMF and torque density,
@@
Henceforth, we will discard the angled brackets and simply bear in mind that we are dealing with the time-averaged results. Integrating these torque densities across the full radial extent and adding the contributions from all modes yields the net torque $T$ enacted on the entire disc. Additionally, the conservation of the planet’s Jacobi integral under the gravitational back-reaction of each uniformly rotating wave mode, relates the planetary orbital energy $E$ to the torque and pattern speed. Indeed one can show that
$\tau_{\rm E}^{-1}\equiv\frac{1}{E}\frac{dE}{dt}=(2/{M_{\rm p}a_{\rm p}^{2}n^{2}})\sum_{m,l}\omega_{ml}T_{ml}.$ (18)
@@
$L=M_{\rm p}\sqrt{GM_{\star}a_{\rm p}(1-e^{2})}=M_{\rm p}na_{\rm p}^{2}\sqrt{1-e^{2}},$ (19)
@@
$E=-\frac{GM_{\star}M_{\rm p}}{2a_{\rm p}}.$ (20)
Differentiating these with respect to time, yields
@@ 2.5 Solution parameters
Equipped with the linear theory and solution methodology summarized above, we are now in a position to outline the target parameter space. Whilst FR25 comprehensively explores a grid of $(q,p,h,e)$ values, here we will be more selective since our emphasis will be placed on detailed comparisons with a discerning number of fully nonlinear, numerical experiments (as described in section 3) which push to higher eccentricities than previously studied. To this end, we will adopt three fiducial disc profiles with $(p,q)\in\{(1.5,1.0),(1.5,0.0),(0.5,0.0)\}$. In each case, we will compute the wake response for a wide range of $e$ ranging between 0.0 and 0.7. For all calculations, we adopt a reference scale height $h_{\rm p}=0.06$ and a smoothing length $\epsilon=0.3h_{\rm p}a_{\rm p}$. Each calculation will adopt a default modal range of $m\in(1,170)$ and $|l-m|\in(0,40)$, which proves to be sufficient for attaining good convergence with the numerical experiments. We address the nature of this convergence with varying $|l-m|$ in appendix A. Note that we are also careful to exclude $(m,l)$ contributions for which the associated pattern speeds yield Lindblad resonances that lie external to the solution domain. In such cases, the attempt to apply radiative wave boundary conditions, when physically there should be no outgoing waves, yields spurious results.
@@ 3 Numerical setup
$\partial_{t}(\Sigma\mathbf{v})+\nabla\cdot(\Sigma\mathbf{v}\mathbf{v}+P\mathbb{I})=-\Sigma\nabla\Phi_{T},\$ (24)
@@
We approximate solutions to Eqns. 2324 using the moving-mesh finite volume code Disco (Duffell, 2016).333 Specifically, we used the version https://github.com/ajdittmann/Disco, which includes additional optimizations and in-situ diagnostics. To achieve very low numerical dissipation when simulating thin (thus having a high azimuthal Mach number) accretion disks, we adopted in Disco a moving-mesh strategy in which each grid annulus rotated with the average angular velocity of the fluid in that annulus. The simulations used second-order Runge-Kutta time stepping (Gottlieb & Shu, 1998) and a second-order total-variation-diminishing spatial scheme (van Leer, 1979; Kurganov & Tadmor, 2000) along with a Harten-Lax-Van Leer-Contact approximate Riemann solver (Toro et al., 1994).
@@
To make direct comparison with the linear calculations, we centre the cylindrical coordinate system of the simulations on the star but neglect the non-inertial forces that should be introduced in such a frame, equivalent to the neglect of the indirect potential in section 2. Thus, the full gravitational potential is given by that of the star in addition to the direct planetary potential introduced in Eqn. 1: $\Phi_{T}=-M_{\star}/r+\Phi$.
@@
[…]
where $\mathbf{R}$ is the position vector of the planet and $\mathbf{v}$ is its velocity. We calculate time-averaged torque density and (wave-driven) angular momentum flux similarly. Defining $\bar{\Sigma}(r)$ as the azimuth and time-averaged surface density and delimited-⟨$\langle$..$\rangle$ … ⟩ as a mass-weighted time- and azimuth average, the average torque density excited in the disk is $dT/dr=2\pi r\bar{\Sigma}\langle\mathbf{r}_{p}\times\nabla\Phi_{T}\rangle$, and the (wave) angular momentum flux $F_{J}=2\pi r^{2}\bar{\Sigma}(\langle v_{r}v_{\phi}\rangle-\langle v_{r}\rangle\langle v_{\phi}\rangle)$.
Our simulations neglect disc self-gravity and are thus largely scale-free (the surface density can be rescaled arbitrarily and we freely chose time units such that $G(M_{p}+M_{\star})=1$). However, unlike the linear theory presented earlier the mass of the secondary necessarily enters the governing equations and controls the strength of various nonlinearities. Apart from a limited number of explorations in section 5.2.2, we set the planetary mass ratio $q_{r}=M_{p}/M_{\star}=10^{-6}$, and thus $M_{p}=0.0046M_{\rm th}$, well within the linear regime.
@@ 4.1 Wake morphology
Figure 2: Maps of the relative surface density perturbation at pericentre for $e=0.06$. The three columns show the disc profiles $(p,q)$ = $(1.5,0.0)$, $(1.5,1.0)$ and $(0.5,0.0)$.
To complement these plots, in Fig. 2 we also examine the surface density maps across the three disc profiles being investigated. For illustration, we adopt the critical value of $e=h_{\rm p}=0.06$ and plot the wake at the pericentre snapshot. The main morphological features are very similar across all disc profiles with only subtle differences. Most notably, in the case with $q=1.0$ the sound speed decreases outwards so the exterior spiral has a smaller pitch angle and is more tightly wound compared with the globally isothermal runs.
It should be acknowledged that the comparison thus far has been fairly qualitative. Whilst one might be tempted to produce difference maps between the simulation and theory results to quantify the discrepancies, these are complicated by the nonlinear damping of the wake structure. Furthermore, slight shifts in the very narrow wake structures can lead to order unity differences. In the upcoming sections we will find that the time/azimuth-averaged and radially integrated diagnostics allow for a more meaningful quantitative comparison.
@@ 4.2 Torque profiles
We will examine in detail the orbit-averaged radial torque density profiles, defined according to equations (14)-(15), for a number of slices through our data. Figs. 3 and 4 show the cases with $(p,q)$ equal to $(1.5,0.0)$ and $(0.5,0.0)$, respectively. In both examples the left-hand column displays the torque density profile across the full numerical domain, whilst the right-hand column zooms in to the fine structure in the inner and outer disc. The results are plotted in characteristic units of $F_{J,0}/a_{\rm p}$ where $F_{J,0}\equiv\Sigma_{\rm p}a_{\rm p}^{4}\Omega_{\rm k,p}^{2}h_{\rm p}^{-3}(M_{\rm p}/M_{\star})^{2}$ is the typical angular momentum exchanged by the interaction. For our choice of code units this simply becomes $F_{J,0}=q_{\rm r}^{2}/h_{\rm p}^{3}$. Each row moving downwards displays the results for increasing values of $e\in\{0.01,0.06,0.12,0.30,0.50\}$, spanning the near-circular regime, via the transonic eccentricity threshold $e=h_{\rm p}$, towards highly eccentric orbits. The solid blue lines denote the simulation results, whilst the dotted yellow line shows the comparison with linear theory. The vertical dashed lines indicate the planetary pericentre and apocentre locations. In all figures, we immediately note the exquisite overlap between the simulations and linear theory, even for the highest values of eccentricity. This simultaneously suggests that the numerical setup is very well behaved and also that the linear theory has included sufficient modes for convergence (see appendix A).
@@ 4.2.1 Behaviour with $e$
As higher values of $e$ are considered, the torque density changes notably. Specifically, when $e=0.06=h_{\rm p}$ the planet performs transonic radial motion relative to the gas, drastically altering the wake structure and associated torques. Similarly to panel (a) in Fig. 5 of FR25 (for $p=q=0.0$), we see that the peaks and troughs develop more radial oscillations in the vicinity of the planet. Pushing into the supersonic regime with $e=0.12$ (the highest value previously considered in FR25) the torque peak and trough ordering has now reversed compared with the circular regime. The location of these maximum and minimum turning points coincide with the apocentre and pericentre of the eccentric planetary orbit (marked by the vertical dashed lines), indicating that most of the angular momentum exchange is facilitated during these orbital phases. As intuitively explained by Papaloizou & Larwood (2000) and Muto et al. (2011), the planet trails/leads the excited wake at apo/pericentre and hence pulls on the gas in the opposite sense to the standard circular planet-disc interaction – thus explaining the reversal in the torque profile. Furthermore, we have previously seen in Fig. 2 that the wake structure in $\delta\Sigma/\Sigma$ is essentially insensitive to the surface density profile, $p$. Therefore, with all else approximately equal, as the surface density becomes shallower, the negative torque contribution on the disc exerted near apocentre becomes stronger whilst the positive torque around pericentre becomes weaker, hence driving a deeper torque reversal (compare Figs. 3 and 4 for $e\geq 0.12$). Additionally, for low $p$ and high $e$, the positive power enacted on the planet at apocentre can eventually dominate and promote outwards orbital migration as we will see later in section 4.4.
@@
Fig. 4, which displays results for $p=0.5$ and $q=0.0$, reinforces the above findings. It is reassuring to see that we are able to attain remarkable agreement with the linear theory across more general disc parameters. The qualitative features are similar, although the magnitude of the torques and the detailed location of some of the peaks have shifted. One might be surprised at the continued good agreement given the fact that corotation torques enter the problem for such general disc profiles. These are known to engage fundamentally nonlinear effects which can change the character of the corotation torque as fluid undergoes ‘horseshoe’ orbits which exchange angular momentum with the planet (e.g. Ward, 1991; Paardekooper & Papaloizou, 2009). However, given the low value of $q_{\rm r}$ it takes a long time for the nonlinear horseshoe effects to manifest and the classical, linear corotation torque is largely preserved at the time of measurement in the simulations (25 orbits). We will discuss the late time behaviour and emergence of the horseshoe drag further in section 5.2.1.
@@ 4.2.2 Torque wiggles
Whilst the torque density is almost reduced to zero in the far field away from the planet, the right hand column panels of Figs. 3 and 4 zoom in to reveal the presence of ‘torque wiggles’. Once again we see remarkable agreement between the numerics and theory emphasising the ability of our simulations to capture subtle features. This phenomenon has previously been observed in circular planet simulations performed by Arzamasskiy et al. (2018) and Dempsey et al. (2020), before Cimerman et al. (2024) provided a detailed analytical understanding. In summary, they arise due to the coupling between the perturbing potential and the net density wave, as it crosses the line intersecting the star and the planet. This relies on the approximately self-similar structure of the density wake and is dominated by the low order $m$ modes as one move radially outwards from the planet.
@@
Figure 5: Diagonal surface density modes $\delta\Sigma_{mm}$ governing the shape of the torque wiggles for the case $(p,q)=(1.5,0.0)$. Panel (a): the Fourier amplitudes for the surface density perturbations normalized by the scaling from the conserved wave action. Panel (b): the phase angle of the modes relative to the dominant spiral arm phase predicted from linear theory. The curves correspond to two values of $e$ at two representative radii. $(e,r/a_{\rm p}):$ (0.0,3.0) = black solid, (0.0,4.0) = red dashed, (0.06,3.0) = orange dashed, (0.06,4.0) = gray solid.
@@
The differing structures exhibited in panel (a) might initially lead one to surmise that the torque density response will be very different. However, at large $r$ the dominant potential component is the low order $\Phi_{11}$ mode which is relatively insensitive to the eccentricity far away from the planet. This then couples to the $\delta\Sigma_{11}$ mode according to equation (15). Fig. 5 clearly shows that these low order modes are almost identical for the two eccentricities considered at large $r$. Therefore the torque wiggles are essentially preserved, even up towards moderate, transonic eccentricities.
However, the profiles observed for the larger eccentricities of $e=0.3$ and $0.5$ begin to exhibit noticeable differences compared with the near circular runs. The high values of $e$, with larger planetary excursions, mean that off-diagonal and higher order modal contributions of the perturbing potential become ever more important in the far-field domain. This breaks the aforementioned reasoning and disturbs the wiggle structure. Presumably far enough away from the planet, the circular perturber wiggle profile can be reinstated, although the feature amplitudes will be very weak by this point.
Finally, although the circular wiggle structure changes in this high $e$ regime, we continue to find good agreement between the simulations and linear framework. Although these methods do begin to present slight discrepancies nearby the planet, there appears to be effective overlap further away, underlining their success at capturing very subtle features.

### v2

@@ 2.4 Important diagnostics
$F_{J}(r)=r^{2}\Sigma(r)\oint\operatorname{Re}[\delta u_{r}(r,\phi)]\operatorname{Re}[\delta u_{\phi}(r,\phi)]d\phi,$ (13)
@@
$\frac{dT}{dr}=-r\oint\operatorname{Re}[\delta\Sigma(r,\phi)]\operatorname{Re}\left[\frac{\partial\Phi}{\partial\phi}\right]d\phi.$ (14)
These quantities can be computed at any instant from the net perturbation maps. Alternatively, pursuing a more elegant approach, one may insert the modal expansion for each perturbed variable, given by equation (3), into equations (13)-(14) and then exploit integral orthogonality conditions to simplify matters (see FR25). Furthermore, we are interested in the orbit-averaged quantities enacted through the operator $\langle\cdot\rangle=(1/T_{\rm p})\oint\cdot dt$, where $T_{\rm p}=2\pi/n_{\rm p}$ is the orbital period of the perturber. This averaging simplifies the expression further and we arrive at equations which encapsulate the individual modal contributions to the AMF and torque density,
@@
Henceforth, we will discard the angled brackets and simply bear in mind that we are dealing with the time-averaged results. Integrating these torque densities across the full radial extent and adding the contributions from all modes yields the net torque $T$ enacted on the entire disc. Additionally, the conservation of the planet’s Jacobi integral under the gravitational back-reaction of each uniformly rotating wave mode, relates the planetary orbital energy $E$ to the torque and pattern speed (see Fairbairn & Rafikov, 2025). Indeed one can show that
$\tau_{\rm E}^{-1}\equiv\frac{1}{E}\frac{dE}{dt}=(2/{M_{\rm p}a_{\rm p}^{2}n_{\rm p}^{2}})\sum_{m,l}\omega_{ml}T_{ml}.$ (18)
@@
$L=M_{\rm p}\sqrt{GM_{\star}a_{\rm p}(1-e^{2})}=M_{\rm p}n_{\rm p}a_{\rm p}^{2}\sqrt{1-e^{2}},$ (19)
@@
$E=-\frac{GM_{\star}M_{\rm p}}{2a_{\rm p}}=-\frac{1}{2}M_{\rm p}n_{\rm p}^{2}a_{\rm p}^{2}.$ (20)
Differentiating these with respect to time (e.g. Ida et al., 2020; Ataiee & Kley, 2021), yields
@@ 2.5 Solution parameters
Equipped with the linear theory and solution methodology summarized above, we are now in a position to outline the target parameter space. Whilst FR25 comprehensively explores a grid of $(q,p,h,e)$ values, here we will be more selective since our emphasis will be placed on detailed comparisons with a discerning number of fully nonlinear, numerical experiments (as described in section 3) which push to higher eccentricities than previously studied. To this end, we will adopt three fiducial disc profiles with $(p,q)\in\{(1.5,1.0),(1.5,0.0),(0.5,0.0)\}$. In each case, we will compute the wake response for a wide range of $e$ ranging between 0.0 and 0.7. For all calculations, we adopt a reference scale height $h_{\rm p}=0.06$ and a smoothing length $\epsilon=0.3h_{\rm p}a_{\rm p}$. Each calculation will adopt a default modal range of $m\in(1,170)$ and $|l-m|\in(0,40)$, which proves to be sufficient for attaining good convergence with the numerical experiments. We address the nature of this convergence with varying |lm||l-m| in appendix A. We are also careful to exclude contributions from certain combinations of $(m,l)$ in our final modal summation. In particular, we ignore modes for which the corresponding pattern speeds resonate with Lindblad locations lying exterior to our solution domain. Otherwise our linear solver tries to impose outgoing radiative boundary conditions when in fact there is no wave excited within the domain, potentially leading to spurious results.
@@ 3 Numerical setup
$\partial_{t}(\Sigma\mathbf{v})+\nabla\cdot(\Sigma\mathbf{v}\mathbf{v}+P\mathbb{I})=-\Sigma\nabla\Phi_{T},\\$ (24)
@@
We approximate solutions to Eqns. 2324 using the moving-mesh finite volume code Disco (Duffell, 2016).444 Specifically, we used the version https://github.com/ajdittmann/Disco, which includes additional optimizations and in-situ diagnostics. To achieve very low numerical dissipation when simulating thin (thus having a high azimuthal Mach number) accretion disks, we adopted in Disco a moving-mesh strategy in which each grid annulus rotated with the average angular velocity of the fluid in that annulus. The simulations used second-order Runge-Kutta time stepping (Gottlieb & Shu, 1998) and a second-order total-variation-diminishing spatial scheme (van Leer, 1979; Kurganov & Tadmor, 2000) along with a Harten-Lax-Van Leer-Contact approximate Riemann solver (Toro et al., 1994).
@@
To make direct comparison with the linear calculations, we centre the cylindrical coordinate system of the simulations on the star but neglect the non-inertial forces that should be introduced in such a frame, equivalent to the neglect of the indirect potential in section 2. For more details on the consistency of this approach, see the recent work by Crida et al. (2025) and Rafikov et al., in prep. Thus, the full gravitational potential is given by that of the star in addition to the direct planetary potential introduced in Eqn. 1: $\Phi_{T}=-M_{\star}/r+\Phi$.
@@
$\left.\frac{dL}{dt}\right|_{i}$ $=\frac{1}{t_{i}-t_{i-1}}\int_{t_{i-1}}^{t_{i}}dt\int\Sigma\mathbf{R}\times\nabla\Phi_{T}dA$ (25)
$\left.\frac{dE}{dt}\right|_{i}$ $=\frac{1}{t_{i}-t_{i-1}}\int_{t_{i-1}}^{t_{i}}dt\int\Sigma\mathbf{v}\cdot\nabla\Phi_{T}dA,$ (26)
where $\mathbf{R}$ is the position vector of the planet and $\mathbf{v}$ is its velocity. We calculate time-averaged torque density and (wave-driven) angular momentum flux similarly. Defining $\bar{\Sigma}(r)$ as the azimuth and time-averaged surface density and $\langle$$\rangle$ is now a mass-weighted time and azimuth average, the average torque density excited in the disk is $dT/dr=2\pi r\bar{\Sigma}\langle\mathbf{r}_{p}\times\nabla\Phi_{T}\rangle$, and the (wave) angular momentum flux $F_{J}=2\pi r^{2}\bar{\Sigma}(\langle v_{r}v_{\phi}\rangle-\langle v_{r}\rangle\langle v_{\phi}\rangle)$.
Our simulations neglect disc self-gravity and are thus largely scale-free (the surface density can be rescaled arbitrarily and we freely chose time units such that $G(M_{p}+M_{\star})=1$). However, unlike the linear theory presented earlier the mass of the secondary necessarily enters the governing equations and controls the strength of various nonlinearities. Apart from a limited number of explorations in section 5.2.2, we set the planetary mass ratio $q_{r}=M_{p}/M_{\star}=10^{-6}$, and thus $M_{p}=0.0046M_{\rm th}$, well within the linear regime. Our fiducial simulations are then run for 25 planetary orbits, which is sufficient time for establishing the linear wake response, whilst not being so long as to start saturating the corotation torque. A selection of longer runs, lasting 200 planetary orbits, will be performed in section 5.2.1 in order to probe this nonlinear saturation effect.
@@ 4.1 Wake morphology
Figure 2: Maps of the relative surface density perturbation at pericentre from our Disco simulations for $e=0.06$. The three columns show the disc profiles $(p,q)$ = $(1.5,0.0)$, $(1.5,1.0)$ and $(0.5,0.0)$.
To complement these plots, in Fig. 2 we also examine the surface density maps from our simulation runs across the three disc profiles being investigated. For illustration, we adopt the critical value of $e=h_{\rm p}=0.06$ and plot the wake at the pericentre snapshot. The main morphological features are very similar across all disc profiles with only subtle differences. Most notably, in the case with $q=1.0$ the sound speed decreases outwards so the exterior spiral has a smaller pitch angle and is more tightly wound whilst the interior wake is less tightly wound, compared with the globally isothermal runs.
Figure 3: Surface density difference map comparing the linear theory and fiducial simulation results for the disc with $(p,q)=(1.5,0.0)$ and $e=0.06$.
It should be acknowledged that the comparison thus far has been fairly qualitative. Whilst one might be tempted to produce difference maps between the simulation and theory results to quantify the discrepancies, these are complicated by the combination of weakly nonlinear, shock-induced damping of the wave and numerical dissipation of the wake structure (discussed further in section 4.5). Furthermore, slight shifts in the very narrow wake structures can lead to order unity differences. Indeed, in Fig. 3 we demonstrate this by plotting the residual surface density differences between the simulations and linear theory for the particular disc profile with $(p,q)=(1.5,0.0)$ and $e=0.06$. We see that there is good agreement in the vicinity of the planet and its coorbital radius, although further away there are clearly growing residuals. Despite these residuals, it should be emphasised that the qualitative wake structure is very good. In the upcoming sections we will find that the time/azimuth-averaged and radially integrated diagnostics allow for a more meaningful quantitative comparison.
@@ 4.2 Torque profiles
We now examine in detail the orbit-averaged radial torque density profiles, defined according to equations (14)-(15), for a number of slices through our data. Figs. 4 and 5 show the cases with $(p,q)$ equal to $(1.5,0.0)$ and $(0.5,0.0)$, respectively. In both examples the left-hand column displays the torque density profile across the full numerical domain, whilst the right-hand column zooms in to the fine structure in the inner and outer disc. The results are plotted in characteristic units of $F_{J,0}/a_{\rm p}$ where $F_{J,0}\equiv\Sigma_{\rm p}a_{\rm p}^{4}n_{\rm p}^{2}h_{\rm p}^{-3}(M_{\rm p}/M_{\star})^{2}$ is the typical angular momentum exchanged by the interaction. For our choice of code units this simply becomes $F_{J,0}=q_{\rm r}^{2}/h_{\rm p}^{3}$. Each row moving downwards displays the results for increasing values of $e\in\{0.01,0.06,0.12,0.30,0.50\}$, spanning the near-circular regime, via the transonic eccentricity threshold $e=h_{\rm p}$, towards highly eccentric orbits. The solid blue lines denote the simulation results, whilst the dotted yellow line shows the comparison with linear theory. The vertical dashed lines indicate the planetary pericentre and apocentre locations. In all figures, we immediately note the exquisite overlap between the simulations and linear theory, even for the highest values of eccentricity. This simultaneously suggests that the numerical setup is very well behaved and also that the linear theory has included sufficient modes for convergence (see appendix A).
@@ 4.2.1 Behaviour with $e$
As higher values of $e$ are considered, the torque density changes notably. Specifically, when $e=0.06=h_{\rm p}$ the planet performs transonic radial motion relative to the gas, drastically altering the wake structure and associated torques. Similarly to panel (a) in Fig. 5 of FR25 (for $p=q=0.0$), we see that the peaks and troughs develop more radial oscillations in the vicinity of the planet. Pushing into the supersonic regime with $e=0.12$ (the highest value previously considered in FR25) the torque peak and trough ordering has now reversed compared with the circular regime. The location of these maximum and minimum turning points coincide with the apocentre and pericentre of the eccentric planetary orbit (marked by the vertical dashed lines), indicating that most of the angular momentum exchange is facilitated during these orbital phases. As intuitively explained by Papaloizou & Larwood (2000) and Muto et al. (2011), the planet trails/leads the excited wake at apo/pericentre and hence pulls on the gas in the opposite sense to the standard circular planet-disc interaction – thus explaining the reversal in the torque profile. Furthermore, we have previously seen in Fig. 2 that the wake structure in $\delta\Sigma/\Sigma$ is essentially insensitive to the surface density profile, $p$. Therefore, with all else approximately equal, as the surface density becomes shallower, the negative torque contribution on the disc exerted near apocentre becomes stronger whilst the positive torque around pericentre becomes weaker, hence driving a deeper torque reversal (compare Figs. 4 and 5 for $e\geq 0.12$). Additionally, for low $p$ and high $e$, the positive power enacted on the planet at apocentre can eventually dominate and promote outwards orbital migration as we will see later in section 4.4.
@@
Fig. 5, which displays results for $p=0.5$ and $q=0.0$, reinforces the above findings. It is reassuring to see that we are able to attain remarkable agreement with the linear theory across more general disc parameters. The qualitative features are similar, although the magnitude of the torques and the detailed location of some of the peaks have shifted. One might be surprised at the continued good agreement given the fact that corotation torques enter the problem for such general disc profiles. In the vicinity of the planet the interaction becomes intrinsically nonlinear which can change the character of the corotation torque as fluid undergoes ‘horseshoe’ orbits which exchange angular momentum with the planet (e.g. Ward, 1991; Paardekooper & Papaloizou, 2009). However, given the low value of $q_{\rm r}$ it takes a long time for the nonlinear horseshoe effects to manifest and the classical, linear corotation torque is largely preserved at the time of measurement in the simulations (25 orbits). We will discuss the late time behaviour and emergence of the horseshoe drag further in section 5.2.1.
@@ 4.2.2 Torque wiggles
Whilst the torque density is almost reduced to zero in the far field away from the planet, the right hand column panels of Figs. 4 and 5 zoom in to reveal the presence of ‘torque wiggles’. Once again we see remarkable agreement between the numerics and theory emphasising the ability of our simulations to capture subtle features. This phenomenon has previously been observed in circular planet simulations performed by Arzamasskiy et al. (2018) and Dempsey et al. (2020), before Cimerman et al. (2024) provided a detailed analytical understanding. In summary, they arise due to the coupling between the perturbing potential and the net density wave, as it crosses the line intersecting the star and the planet. This relies on the approximately self-similar structure of the density wake and is dominated by the low order $m$ modes as one move radially outwards from the planet.
@@
Figure 6: Diagonal surface density modes $\delta\Sigma_{mm}$ governing the shape of the torque wiggles for the case $(p,q)=(1.5,0.0)$. Panel (a): the Fourier amplitudes for the surface density perturbations normalized by the scaling from the conserved wave action. Panel (b): the phase angle of the modes relative to the dominant spiral arm phase predicted from linear theory. The curves correspond to two values of $e$ at two representative radii. $(e,r/a_{\rm p}):$ (0.0,3.0) = black solid, (0.0,4.0) = red dashed, (0.06,3.0) = orange dashed, (0.06,4.0) = gray solid. The eccentric case exhibits an oscillation with nodal spacings indicated by the dotted blue lines.
@@
The differing structures exhibited in panel (a) might initially lead one to surmise that the torque density response will be very different for the circular e=0.0 case versus the eccentric e=0.06 case. However, at large $r$ the dominant potential component is the low order $\Phi_{11}$ mode which is relatively insensitive to the eccentricity far away from the planet. This then couples to the $\delta\Sigma_{11}$ mode according to equation (15). Fig. 6 clearly shows that these low order modes are almost identical for the two eccentricities considered at large $r$. Therefore the torque wiggles are essentially preserved, even up towards moderate, transonic eccentricities. Indeed, whilst we have not plotted the corresponding $e=0.12$ curves in Fig. 6 so as to maintain clarity, they essentially paint the same picture wherein there is close overlap with the $e=0.0$ curves for the lowest order modes.
However, the profiles observed for the larger eccentricities of $e=0.3$ and $0.5$ begin to exhibit noticeable differences compared with the near circular runs. The high values of $e$, with larger planetary excursions, mean that off-diagonal and higher order modal contributions of the perturbing potential become ever more important in the far-field domain. Furthermore, the low order, diagonal density perturbations also begin deviating from the lower eccentricity curves plotted in Fig. 6. Together, this breaks the aforementioned reasoning and disturbs the wiggle structure. Presumably far enough away from the planet, the circular perturber wiggle profile can be reinstated, although the feature amplitudes will be very weak by this point.
Finally, although the circular wiggle structure changes in this high $e$ regime, we continue to find good agreement between the simulations and linear framework. Although these methods do begin to present slight discrepancies near the planet, there appears to be effective overlap further away, underlining their success at capturing very subtle features.